# Building spatial world models from sparse transitional episodic memories

### v1

@@ Abstract
Many animals possess a remarkable capacity to rapidly construct flexible mental models of their environments. These world models are crucial for ethologically relevant behaviors such as navigation, exploration, and planning. The ability to form episodic memories and make inferences based on these sparse experiences is believed to underpin the efficiency and adaptability of these models in the brain. Here, we ask: Can a neural network learn to construct a spatial model of its surroundings from sparse and disjoint episodic memories? We formulate the problem in a simulated world and propose a novel framework, the Episodic Spatial World Model (ESWM), as a potential answer. We show that ESWM is highly sample-efficient, requiring minimal observations to construct a robust representation of the environment. It is also inherently adaptive, allowing for rapid updates when the environment changes. In addition, we demonstrate that ESWM readily enables near-optimal strategies for exploring novel environments and navigating between arbitrary points, all without the need for additional training.
@@ 1 Introduction
When visiting a new city, we quickly piece together scattered memories—an alley passed on a walk, a glimpse of a landmark from a taxi—into a mental map that lets us navigate, explore, and plan. This ability to build a coherent model of the world from fragmented experiences, and to use it flexibly for imagining new routes or outcomes, is a core feature of human intelligence that is fundamental to planning, problem-solving, and decision-making, all of which are crucial for survival (4).
@@
How do humans achieve this extraordinary feat? Insights from neuroscience and psychology provide compelling evidence for the neural mechanisms underpinning this capacity. One pivotal discovery is the dual role of a key brain area called the medial temporal lobe (MTL) in both spatial representations and episodic memories. Individual neurons within the MTL exhibit spatial selectivity, firing in response to specific locations (39; 23), while collectively forming a population code that represents an animal’s instantaneous position in space (59). In parallel, MTL plays a critical role in the formation of episodic memories, enabling the rapid acquisition of associations between arbitrary stimuli (27). Strikingly, lesions to this area result in profound deficits: individuals lose not only their ability to explore and navigate effectively (32) but also their capacity to form episodic memories (45) and imagine novel scenarios (25). These converging findings strongly suggest that the MTL constructs structured relational networks, integrating overlapping episodic memories into a cohesive framework (12; 35). This mechanism may underlie a flexible and adaptive world model, allowing for inferences about spatial relationships beyond immediate perception.
@@
Despite these insights, most prior computational approaches to world modeling have relied on learning from continuous sequences of recent observations and actions(19; 20; 54; 34), fixed hand-designed circuitry for representing and updating spatial information (52; 5; 33), allocentric observations (30), and iterative update procedures that hinders their ability to quickly adapt to environmental changes(52; 5; 47). These models are typically trained on large datasets collected from a given environment, with knowledge encoded in their weights. While effective in certain contexts, such approaches face inherent limitations. For instance, agents may encounter different parts of an environment at widely separated times, visit states along arbitrary and unrelated trajectories that do not form continuous sequences, or experience environments that undergo structural changes (Fig.1a). These challenges hinder the ability of sequence-based models to generalize effectively in environments where observations are disjoint and episodic.
@@
This raises a key question: Can a neural network rapidly construct a coherent spatial map of the environment using only sparse and disjoint episodic memories? Drawing inspiration from the neuroscience of memory and navigation, we introduce the Episodic Spatial World Model (ESWM), a framework designed to infer spatial structure from independently acquired episodic experiences. While in conventional world-model training, a network’s weights encode the full dynamics of a single environment via long, continuous trajectories, ESWM meta-trains the model to predict unseen transitions from a minimal set of one-step transitions sampled across diverse environments. This approach allows the model to infer missing information, connect distant experiences, and dynamically adapt to structural changes.
Our main contributions are as follows:
  1. 1.
    We introduce ESWM, a framework for modeling spatial environments from a minimal set of previously experienced episodic transitions in each environment.
  2. 2.
    We demonstrate that the ESWM’s latent space organizes into a geometric map that closely reflects the spatial structure of the environment.
  3. 3.
    By studying the behavior of the model across various scenarios, we show that ESWM’s latent structure dynamically updates in response to changes in its memories, enabling it to incorporate previously unseen regions and adapt to structural modifications, such as the addition of obstacles.
  4. 4.
    We demonstrate that the learned world model enables near-optimal exploration of novel environments and facilitates navigation between arbitrary points within those environments—all without requiring additional training.
@@
Figure 1: Episodic Spatial World Model. a) Three common scenarios that hinder the ability of typical world models to generalize effectively. (top) Observation of the full environment may take many time steps leading to long sequences; (middle) Specific parts of the environment may be changed dynamically; (bottom) Information about a particular environment may be collected across separate exposures to the environment and not within a single one. b) . Memory bank and query selection in a square grid environment. c) Architecture of ESWM and training procedure. Model input consists of a bank of transitional memories (corresponding to the black arrows in (b)) and a single query (q arrow in (b)), with either start-state, action, or end-state randomly masked with equal chances. The sequence of transitions is processed by a sequence model (e.g, Transformer encoder block) and the model parameters are updated to output the correct value for the masked component.
@@ 2.1 World models
Humans excel at forming abstract representations of spatial and temporal information from egocentric observations and action, enabling efficient interaction with complex, dynamic environments. Inspired by this cognitive ability, prior work in artificial intelligence, particularly in reinforcement learning (RL) (48), has explored world models to improve sample efficiency and planning effectiveness (22; 44; 50). World models provide agents with a predictive understanding of their environment by modeling past and present states and forecasting future outcomes (19). In spatial settings, agents have been trained to predict future sensory observations following arbitrary actions (47; 54; 34) while in vision, methods such as image reconstruction (53; 1) and video prediction (28; 18) enable simulation of future frames or latent dynamics. These approaches have demonstrated success in domains such as video games (21; 44; 28) and robotic control tasks (18; 11).
More recently, the emergence of transformer architectures (51) has opened new possibilities for improving world models. Recent works such as TWM (43), STORM (61), and IRIS (36) have leveraged transformers-based world models to improve performance on sample-limited benchmarks like Atari 100k. However, unlike our proposed model, these models train extensively on a single environment and absorb the knowledge about that environment in the model’s weight.
@@ 2.2 Navigation Agents
The concept of a spatial or cognitive map (40) has been thought to enable navigation and spatial reasoning. This spatial map, an internal neural representation of an environment, encodes landmarks, resources, and shelters, enabling organisms to plan and navigate (39; 13; 56). Several studies have demonstrated that similar spatial representations emerge in neural networks trained on navigation tasks (16). For example, Banino et al. (2) and Cueva and Wei (10) showed that spatial representations arise in recurrent neural networks trained for path integration, aiding navigation through implicit spatial reasoning. In parallel, neural networks lacking explicit mapping modules have achieved high performance on navigation tasks, demonstrating that general-purpose architectures can achieve high navigation accuracy in novel environments (57; 41; 31; 46). For instance, Reed et al. (41) showed that a general-purpose transformer model could perform a wide range of embodied tasks, including navigation.
Neuroscience findings suggest that the brain employs common cognitive mapping strategies across both spatial and non-spatial domains, potentially encoding abstract information like semantic content within these maps (3). For instance, this perspective frames reasoning as a form of navigation through a cognitive map of concepts, relationships, and logical rules. Wijmans et al. (58) recently showed that spatial representations can emerge in navigation-agnostic neural network architectures trained for navigation, emphasizing memory’s role in shaping these spatial representations. This raises a key question: Can spatial maps also emerge in general-purpose models not explicitly trained for navigation?
Figure 2: Evaluation accuracy in a) Open Arena and b) Random Wall. In both schematics, blue arrows are transitions in the memory bank and the $q$-labeled arrows are examples of query transitions $q=(s_{s},a,s_{e})$. In Open Arena, the states are represented as 6-bit binaries and the model is tested in environments entirely filled with unseen states, while in Random Wall, the states are integers and the model is tested on its generalizability to unseen wall patterns. In addition, a random subregion is masked in Random Wall and the query $q$ can be either unsolvable, unseen, or seen, depicted as $q$-labeled arrows with different colors. The Random Wall displayed here is a scaled down version of what the model is trained on (19 locations versus 36 locations). XFMR is short for Transformer.
@@ 3.1 Episodic Spatial World Model
We define an episodic memory as a one-step transition in a spatial environment, represented by the tuple $(s_{s},a,s_{e})$—denoting the source state, action, and end state, respectively. A memory bank $M$ for a given environment is an unordered set of such transitions, constructed to be: Disjoint, transitions do not form a continuous trajectory; Spanning, the transitions collectively form a connected graph covering all locations; Minimal, removing any transition would disconnect the graph (Fig. 1b). We algorithmically generate such memory banks to sparsely cover each environment (Extended Methods A.3). Importantly, a single room may yield multiple valid memory banks, enabling our meta-training setup.
We formulate spatial modeling as the ability to infer any missing component of an unseen but plausible transition $q$—that is, a query transition not present in $M$ (Fig.1b). The task is to predict the masked element (either $s_{s}$, $a$, or $s_{e}$) of $q$, given the other two and the memory bank $M$ (Fig.1c).
@@
To train ESWM, we adopt a meta-learning approach where on each trial (i.e. sample), we randomly sample: (1) An environment $e$; (2) A disjoint memory bank $M$ from $e$; (3) A plausible query transition $q$ from ($e\setminus M$); (4) A masking choice (randomly masking $s_{s}$, $a$, or $s_{e}$). This process ensures that the model must reason over episodic structure rather than memorize specific environments. Notably, our environments (Open Arena, Random Wall; detailed below) contain up to $10^{33}$ and $5\times 10^{36}$ possible configurations, respectively—making memorization intractable (Extended Methods A.2, Fig. 2).
@@ 3.2 Model Selection and Training
We considered three neural network architectures for $f$, 1) Encoder-only transformer with no positional encoding(51), 2) long short-term memory (LSTM) (26), and Mamba (17). To create an embedding of each transition $(s_{\text{s}},a,s_{\text{e}})$, each component of the transition (i.e. source state, action, and end state) are projected onto the same dimensional space using separate linear projections. The three projected values are averaged to produce a single embedding of the transition.
Each model receives as input the concatenated sequence of randomly-permuted memory bank embeddings and query embeddings, treating each embedding (i.e. each transition) as a distinct input token (Fig.1b). The LSTM and Mamba model receives a sequence of tokens that start with $M$ and end with the query token. Three separate linear heads read out the model’s prediction on the three parts $(s_{\text{s}},a,s_{\text{e}})_{q}$ from the final layer’s last token. We use cross-entropy loss between predicted and actual states and actions. Equal weight is assigned to each prediction ($s_{s},a,s_{e}$). Training is done over 460k iterations with a batch size of 128 and a cosine learning rate schedule.
@@ 3.3 Model Evaluation
We evaluate ESWM in a family of discrete, hexagonally structured grid environments. The hexagonal layout allows for six possible actions per location, offering richer connectivity than traditional square grids. Each location in the environment is assigned an integer-valued observation, drawn from a predefined set of states. An environment is uniquely defined by its mapping from states to spatial locations, allowing us to generate a vast number of distinct rooms by randomly shuffling this mapping—critical for our meta-learning objective (Extended Methods A.2).
We design two types of environments to test generalization: (1) Open Arena to evaluate generalization to unseen observations and; (2) Random Wall to evaluate generalization to unseen structures.
@@
Open Arena: Each room consists of 19 locations and 64 possible state values. These states are represented compositionally as 6 binary bits (Fig. 2a). For action prediction, the model performs 6-way classification (one for each action). For source and end state prediction, the task is broken into six independent binary classification problems—one for each bit. To test generalization, we ensure that training and testing states are mutually exclusive: the model is never exposed to the same state values during training and testing. This forces it to build representations that support spatial reasoning even in environments filled with previously unseen but compositionally generated state observations.
Random Wall: In this setting, rooms are larger (36 locations) and contain 36 possible states, each represented as a unique integer. The association between locations and states is still randomized for each sample. Importantly, each room also includes a randomly shaped wall composed of obstacles that block movement (i.e., attempting to enter a wall cell returns the agent to its original location) (Fig. 2b). The model must generalize to unseen state-location associations and wall configurations during testing. Prediction tasks mirror the Open Arena setting: 6-way classification for actions, and 36-way classification for source and end states.
We also allow the model to output “I don’t know” (implemented as an additional classification category) when the prediction is not possible given the available memory bank. To simulate such scenarios, we occasionally create trials in which some of the memories covering a part of the environment are removed from the memory bank and query the model with a transition that crosses between known and unknown regions (“unsolvable queries”). This enables the model to identify uncertainty and encourages adaptive exploration in novel, obstacle-rich environments. During training, query transitions are sampled with the following probabilities: 68% unseen (plausible but not in the memory bank), 17% seen (present in the memory bank), 15% unsolvable (crossing into unknown regions) (Appendix A.4).

### v2

@@ Abstract
Many animals possess a remarkable capacity to rapidly construct flexible cognitive maps of their environments. These maps are crucial for ethologically relevant behaviors such as navigation, exploration, and planning. Existing computational models typically require long sequential trajectories to build accurate maps, but neuroscience evidence suggests maps can also arise from integrating disjoint experiences governed by consistent spatial rules. We introduce the Episodic Spatial World Model (ESWM), a novel framework that constructs spatial maps from sparse, disjoint episodic memories. Across environments of varying complexity, ESWM predicts unobserved transitions from minimal experience, and the geometry of its latent space aligns with that of the environment. Because it operates on episodic memories that can be independently stored and updated, ESWM is inherently adaptive, enabling rapid adjustment to environmental changes. Furthermore, we demonstrate that ESWM readily enables near-optimal strategies for exploring novel environments and navigating between arbitrary points, all without the need for additional training. Our work demonstrates how neuroscience-inspired principles of episodic memory can advance the development of more flexible and generalizable world models.
@@ 1 Introduction
When visiting a new city, we quickly piece together scattered memories—an alley passed on a walk, a glimpse of a landmark from a taxi—into a mental map that lets us navigate, explore, and plan. This ability to build a coherent model of the world from fragmented experiences, and to use it flexibly for imagining new routes or outcomes, is a core feature of human intelligence that is fundamental to planning, problem-solving, and decision-making, all of which are crucial for survival (Bennett, 2023).
@@
How do humans achieve this extraordinary feat? Insights from neuroscience and psychology provide compelling evidence for the neural mechanisms underpinning this capacity. One pivotal discovery is the dual role of a key brain area called the medial temporal lobe (MTL) in both spatial representations and episodic memories. Individual neurons within the MTL exhibit spatial selectivity, firing in response to specific locations (O’Keefe, 1976; Hafting et al., 2005), while collectively forming a population code that represents an animal’s instantaneous position in space (Wilson & McNaughton, 1993). In parallel, MTL plays a critical role in the formation of episodic memories, enabling the rapid acquisition of associations between arbitrary stimuli (Howard & Eichenbaum, 2015). Strikingly, lesions to this area result in profound deficits: individuals lose not only their ability to explore and navigate effectively (Kimble, 1963) but also their capacity to form episodic memories (Scoville & Milner, 1957) and imagine novel scenarios (Hassabis et al., 2007). These converging findings strongly suggest that the MTL constructs structured relational networks, integrating overlapping episodic memories into a cohesive framework (Eichenbaum, 2004; McKenzie et al., 2014) (See Appendix B for more a detailed review). This mechanism may underlie a flexible and adaptive world model, allowing for inferences about spatial relationships beyond immediate perception.
@@
Despite these insights, most prior computational approaches to world modeling have relied on learning from continuous sequences of recent observations and actions (Ha & Schmidhuber, 2018; Hafner et al., 2019; Whittington et al., 2020; Levenstein et al., 2024; Bar et al., 2025; Fraccaro et al., 2018), fixed hand-designed circuitry for representing and updating spatial information (Wang et al., 2023; Chandra et al., 2025; Kymn et al., 2024), allocentric observations Khan et al. (2018), and iterative update procedures that hinders their ability to quickly adapt to environmental changes(Wang et al., 2023; Chandra et al., 2025; Stachenfeld et al., 2017). These models are typically trained on large datasets collected from a given environment, with knowledge encoded in their weights. While effective in certain contexts, such approaches face inherent limitations. For instance, agents may encounter different parts of an environment at widely separated times, visit states along arbitrary and unrelated trajectories that do not form continuous sequences, or experience environments that undergo structural changes (Fig. 1a). These challenges hinder the ability of sequence-based models to generalize effectively in environments where observations are disjoint and episodic.
@@
This raises a key question: Can a neural network (NN) rapidly construct a coherent spatial map of the environment using only sparse and disjoint episodic memories? Drawing inspiration from the neuroscience of memory and navigation, we introduce the Episodic Spatial World Model (ESWM), a framework designed to infer spatial structure from independently acquired episodic experiences. ESWM meta-trains the model to predict unseen transitions from a sparse set of one-step transitions (referred to as the memory bank) sampled across diverse environments. This is in contrast with conventional world models in two ways: 1) ESWM does not require sequential observations for predictions–the memory bank can contain transitions from completely different episodes. This is crucial in large environments where a single walk spanning all locations might be long and therefore computationally expensive to process. In comparison, ESWM’s memory bank can be constructed to contain only the useful transitions. 2) Each transition in the memory bank can be stored and updated independently. As a result, ESWM can rapidly adapt to environmental changes (e.g, addition of obstacles) by simply editing a handful of transitions that are related to that change (see Section 4.6).
Our main contributions are: 1) We introduce ESWM, a framework that models spatial environments from sparse, episodic transitions. 2) We find that ESWM’s latent space forms a geometric map reflecting the environment’s topology, which dynamically adapts to new memories and structural changes like obstacles. 3) We demonstrate that this learned model enables near-optimal, zero-shot exploration and navigation in novel environments without any task-specific training. 4) Finally, we show that ESWM is scalable to realistic continuous environments with high-dimensional inputs.
@@
Figure 1: Episodic Spatial World Model. a) Three common scenarios that hinder the ability of typical world models to generalize effectively. (top) Observation of the full environment may take many time steps leading to long sequences; (middle) Environment structure may change across trials; (bottom) Information about a particular environment may be collected across separate exposures to the environment and not within a single one. b) Memory bank and query selection in a square grid environment. c) Architecture of ESWM and training procedure. Model input consists of a bank of transitional memories (corresponding to the black arrows in (b)) and a single query (q arrow in (b)), with either start-state, action, or end-state randomly masked with equal chances. The sequence of transitions is processed by a sequence model (e.g, Transformer encoder block) and the model parameters are updated to output the correct value for the masked component.
@@ 2.1 World models
Humans excel at forming mental models of the world from egocentric observations and action, enabling efficient interaction with complex, dynamic environments. Inspired by this cognitive ability, prior works have used NN to forecast future outcomes based on past and present states (Ha & Schmidhuber, 2018) using methods such as image reconstruction (Watter et al., 2015; Alonso et al., 2024) in domains like video games (Hafner et al., 2020; Schrittwieser et al., 2020; Kaiser et al., 2019) and robotic control tasks (Gupta et al., 2022; Ebert et al., 2018). In spatial settings, agents have been trained to predict future sensory observations following arbitrary actions, typically using recurrently-connected networks (Stachenfeld et al., 2017; Whittington et al., 2020; Levenstein et al., 2024; Fraccaro et al., 2018).
More recently, the emergence of transformer architectures (Vaswani et al., 2017) has opened new possibilities for improving world models. Recent works such as TWM (Robine et al., 2023), STORM (Zhang et al., 2024), and IRIS (Micheli et al., 2022) have leveraged transformers-based world models to improve performance on sample-limited benchmarks like Atari 100k. However, unlike our proposed model, these models train extensively on a single environment and absorb the knowledge about that environment in the model’s weight.
Most notably, our work is closely related to Generative Temporal Models with Spatial Memory (GTM-SM) Fraccaro et al. (2018) and the Tolman-Eichenbaum Machine (TEM) (Whittington et al., 2020; 2022), the latter of which has been proposed as a model of medial temporal lobe function. These models factorize sensory and structural information. The shared structure is encoded in a recurrent network weights, while environment-specific observations are stored in Hebbian-slot-memory, self-attention module, or differentiable memory dictionary. While both of these model classes include content addressable memory modules, important distinctions separate them from our proposed model. First, GTM-SM and TEM both presuppose a common structural template across environments (e.g, an open 2D grid), whereas ESWM does not. Second, unlike these models, ESWM does not embed structural knowledge in its weights but instead infers it dynamically from external memories, enabling it to capture distributions of environments with varying structures (e.g, a set of 2D mazes) where these models fail (Fig. 2). Finally, while GTM-SM and TEM operate over sequential trajectories, ESWM integrates disjoint experiences, making it more sample-efficient and better suited to rapid adaptation, such as handling newly introduced obstacles. See Appendix A for a more comprehensive comparison with other MTL models.
@@ 2.2 Navigation Agents
The concept of cognitive map, an internal allocentric representation of the space, O’keefe & Nadel (1978) has been thought to enable navigation and spatial reasoning. Several studies have demonstrated that similar spatial representations emerge in NN trained on navigation tasks (Gornet & Thomson, 2024). For example, Banino et al. (2018) and Cueva & Wei (2018) showed that spatial representations arise in recurrent neural networks trained for path integration, while in parallel, NN lacking explicit mapping modules have achieved high performance on navigation tasks in unseen environments (Wijmans et al., 2019; Reed et al., 2022; Khandelwal et al., 2022; Shah et al., 2023). Wijmans et al. (2023) recently showed that spatial representations can emerge in navigation-agnostic NN architectures trained for navigation, emphasizing memory’s role in shaping these spatial representations. This raises a key question: Can spatial maps also emerge in general-purpose models not explicitly trained for navigation and operating entirely on a finite set of episodic memories?
@@ 3.1 Episodic Spatial World Model
We define an episodic memory as a one-step transition in a spatial environment, represented by the tuple $(s_{s},a,s_{e})$—denoting the source state, action, and end state, respectively. For each environment, we construct a memory bank $M$ as an unordered set of such transitions, constructed to be: Disjoint, transitions do not form a continuous trajectory; Spanning, the transitions collectively form a connected graph covering all locations; Minimal, removing any transition would disconnect the graph (Fig. 1b). However, Minimality is not required for training ESWM but is imposed in the grid environment to prove that the absolute minimum information is sufficient to construct a functional spatial map (see Fig. 9 for results on non-minimal memory banks). We algorithmically generate such memory banks to sparsely cover each environment (Extended Methods C.3). Importantly, a single room may yield multiple valid memory banks, enabling our meta-training setup.
We formulate spatial modeling as the ability to infer any missing component of an unseen but plausible transition qq—that is, a query transition not present in $M$ (Fig. 1b). The task is to predict the masked element (either $s_{s}$, $a$, or $s_{e}$) of $q$, given the other two and the memory bank $M$ (Fig. 1c). For example: predicting the next state from $s_{s}=5$ and $a=\text{move right}$; inferring the action connecting states 5 and 10; or recovering the start state given $a=\text{move right}$ and $s_{e}=10$.
To train ESWM, we adopt a meta-learning approach where on each trial (i.e. sample), we randomly sample: (1) An environment $e$ (though random assignment of states to locations); (2) A disjoint memory bank $M$ from $e$; (3) A plausible query transition $q$ from ($e\setminus M$); (4) A masking choice (randomly masking $s_{s}$, $a$, or $s_{e}$). This process prevents the model from memorizing any specific environment since the state-location mapping is randomized across trials (i.e, the model cannot simply memorize that state ”5” is at top-left because ”5” could be at top-right the next trial), and our environments (Open Arena, Random Wall; detailed below) contain an intractable amount of configurations ($10^{33}$, see Extended Methods C.2, Fig. 2).
@@ 3.2 Model Selection and Training
We considered three NN architectures for $f$, 1) Encoder-only transformer with no positional encoding(Vaswani et al., 2017), ESWM-T, 2) long short-term memory (LSTM) (Hochreiter & Schmidhuber, 1997), ESWM-LSTM, and Mamba (Gu & Dao, 2024), ESWM-MAMBA. Each transition component $(s_{s},a,s_{e})$ is projected separately into a shared high-dimensional space and averaged to form a single input token.
The transformer model receives as input the concatenated sequence of memory bank and query tokens (Fig. 1b). The LSTM and Mamba models receive a randomly permuted sequence of memory tokens, each with the query token added. Three separate linear heads read out the model’s prediction on the three parts $(s_{\text{s}},a,s_{\text{e}})_{q}$ from the final layer’s last token. We use cross-entropy loss between predicted and actual states and actions. Equal weight is assigned to each prediction ($s_{s},a,s_{e}$). Training is done over 460k iterations with a batch size of 128 and a cosine learning rate schedule.
A spatial world model that assumes complete memory coverage is rather constrained. Instead, it should reason over partial observations. To achieve this, we create trials where some memories are deliberately removed. The model is expected to solve queries within observed regions and classify unsolvable ones (e.g., transitions involving unobserved regions) under an additional “I don’t know” category (Fig. 2b).
@@ 3.3 Model Evaluation
Figure 2: Evaluation accuracy in a) Open Arena and b) Random Wall. In both schematics, blue arrows are transitions in the memory bank and the $q$-labeled arrows are examples of query transitions $q=(s_{s},a,s_{e})$. In Open Arena, the states are represented as 6-bit binaries and the model is tested in environments entirely filled with unseen states, while in Random Wall, the states are integers and the model is tested on its generalizability to unseen wall patterns. In addition, a random subregion is masked in Random Wall and the query $q$ can be either unsolvable, unseen, or seen, depicted as $q$-labeled arrows with different colors. The Random Wall displayed here is a scaled down version of what the model is trained on (19 v.s 36 locations).
We evaluate ESWM under three settings of varying complexity:
  1. 1.
@@
    2D grid environments: We evaluate ESWM in a family of discrete, hexagonally structured grid environments (Extended MethodsC.2). Each location in the grid environment is assigned an integer-valued observation, drawn from a predefined set of states. An environment is uniquely defined by its mapping from states to spatial locations, allowing us to generate a vast number of distinct rooms by randomly shuffling this mapping—critical for our meta-learning objective. The hexagonal layout allows for six possible actions per location, offering richer connectivity than traditional square grids. We design two types of environments to test generalization: (1) Open Arena to evaluate generalization to unseen states and; (2) Random Wall to evaluate generalization to unseen structures.
@@
    • Open Arena: Each room consists of 19 locations and 64 possible state values. These states are represented compositionally as 6 binary bits (Fig. 2a), enabling us to train and test ESWM on mutually exclusive sets of states (e.g, train on odd states and test on even states). For action prediction, the model performs 6-way classification (one for each action). For source and end state prediction, the task is broken into six independent binary classification problems—one for each bit. This setup forces it to build representations that support spatial reasoning even in environments filled with previously unseen but compositionally generated state observations.
    • Random Wall: To test for generalization to unseen structures, this environment features larger rooms (36 locations) with unique integer states and a randomly shaped wall of obstacles (Fig. 2b). During testing, the model must generalize to novel wall configurations and state-location mappings. The prediction task is 6-way classification for actions and 36-way for states. We simplified the observations to integers to focus the task on inferring the environment’s underlying structures.
  2. 2.
    MiniGrid environments: We used MiniGrid (Chevalier-Boisvert et al., 2023) to create environment layouts of size $9\times 9$ with egocentric multi-dimensional observations and actions (Fig. 6). Each state consist of an egocentric $5\times 5$ view of the environment containing non-unique colored objects. The agent takes one of three egocentric actions: turn left, turn right, go straight. Each procedurally generated environment is assigned a random color layout and wall pattern. The model has to self-localize based on partial egocentric views of the environment experienced from varying head directions.
  3. 3.
    Simulated 3D indoor environments: We use ProcThor (Deitke et al., 2022), a large set of procedurally generated 3D indoor scenes. Each state consists of a high-resolution RGB first-person observation ($224\times 224$). The agent takes action defined by egocentric displacements $\Delta xy$ (along left/right and front/back axes) and rotation $\Delta\theta$, both from a continuous range.
For MiniGrid and ProcThor environments, the memory bank consists of episodic memories that collectively cover the room, with partial overlap across observations to enable memory integration. Importantly, the memory banks in these environments remain compact.